# Cortex AISQL: A Production SQL Engine for Unstructured Data

### v2

@@ 5.2. Adaptive Model Cascades
AI operators such as AI_FILTER must often process millions of rows in production database workloads. However, invoking an LLM on every row introduces prohibitive costs and latency. To address this challenge, the AISQL engine implements model cascades: a lightweight proxy model processes most rows, while a powerful oracle model handles only difficult cases where the proxy exhibits uncertainty. The proxy model’s predictions are accepted when confidence is high, but uncertain cases are escalated to the oracle for more reliable decisions.
Implementing model cascades in a production database system presents unique challenges. First, the system must learn an effective routing strategy online from streaming data without prior knowledge of the query’s data distribution. Second, the algorithm must operate in a highly concurrent environment, where multiple workers process different portions of the data in parallel. Third, it must honor user-specified cost budgets while maintaining high-quality results. Finally, the routing decisions must adapt dynamically as data characteristics change across batches.
Snowflake’s AISQL engine addresses these challenges through an adaptive threshold learning algorithm that reframes cascade routing as a statistical estimation problem. The algorithm uses a sample of oracle evaluations to learn confidence thresholds that partition rows into three regions: an accept region where the proxy is confident and predicts positive, a reject region where proxy confidence is low and predicts negative, and an uncertainty region that might require invoking the oracle and making conservative decisions.
The algorithm operates in four phases over batches of rows:
Phase 1: Proxy Inference. Execute the proxy model on all rows in the current batch to obtain predictions and confidence scores. For an AI_FILTER predicate, each row receives a score representing the proxy’s confidence that the predicate is satisfied.
Phase 2: Sampling. Allocate a fraction of the oracle budget to sample rows for ground-truth labels. Rather than sampling uniformly, the algorithm uses importance sampling to oversample rows with intermediate confidence scores. These are the rows where the proxy is most uncertain and where oracle feedback provides maximum information for threshold estimation.
Phase 3: Threshold Learning. Using the sampled oracle labels, compute two thresholds and that define the uncertainty region:
Rows with are predicted negative (proxy is confident it’s false)
Rows with are predicted positive (proxy is confident it’s true)
Rows with remain in the uncertainty region
These thresholds are computed using concentration inequalities that provide probabilistic guarantees with limited samples (kang2020approximate). The algorithm accounts for sampling bias through importance weighting, ensuring threshold estimates are unbiased despite non-uniform sampling.
Phase 4: Adaptive Refinement. The algorithm refines thresholds iteratively across batches. For each new batch: (1) identify rows falling in the current uncertainty region, (2) allocate remaining oracle budget to sample additional rows from this region, (3) update threshold estimates using the new sample, and (4) repeat until the oracle budget is exhausted or the uncertainty region is empty. Iterative refinement allows the thresholds to adapt to evolving data distributions.
Once thresholds are learned, the engine makes final routing decisions. Rows outside the uncertainty region are classified directly using learned thresholds and proxy predictions. Rows remaining in the uncertainty region are routed to the oracle if budget permits; otherwise, the proxy prediction is used as a fallback.
The iterative batch-based implementation in the AISQL engine naturally accommodates concurrent database execution. Multiple workers process different data partitions in parallel, each collecting oracle samples within their assigned batches. These samples may be periodically aggregated to update shared threshold estimates, allowing the algorithm to learn from the entire data stream while workers proceed independently.
@@ 5.3. Query Rewriting for Semantic Joins
Semantic joins using AI_FILTER enable powerful natural language-based table joining capabilities. For example, let us consider a scenario where a user has two Snowflake tables: (1) “Reviews” contains product reviews, and (2) “Categories” includes names of product categories. A simplified schema of these two tables, along with some sample data, is shown in Figure 8.
@@
Due to the semantic nature of the join condition, efficient join algorithms such as hash-join or sort-merge join, which rely on equality predicates, are not applicable. The AISQL execution engine must therefore treat the semantic join as a cross join, thereby generating all possible pairs from Reviews and Categories and evaluating AI_FILTER for each pair. Therefore, many queries like this can result in thousands of LLM calls and high inference costs, even for small datasets. More specifically, the above approach requires LLM calls, where and denote the cardinalities of the respective tables.
After analyzing a wide range of customer queries, we discovered that, in many cases, semantic joins are equivalent to a multi-label classification problem. For instance, in the previous example, rather than applying AI_FILTER for every review-category pair, we can leverage the Snowflake AI_CLASSIFY function, which can classify a text or image input into one or more user-specified labels. By re-framing the semantic join into a classification problem we can instead execute AI_CLASSIFY for every row of table Reviews using the values of column Categories.label as labels, thereby reducing the number of AI inference calls from 24 (4 reviews 6 labels) to just 4 (one for each review), a 6 improvement in the example above.
Automatically detecting such cases is not a trivial task. During the compilation of AISQL queries, we introduced an AI-based oracle that examines every semantic join to determine if it can be rewritten as a multi-label classification problem. The oracle analyzes the user-specified natural language prompt, schema metadata (e.g., column and table names), statistics (e.g., number of distinct values), as well as sample values from each input source. Using all that information, the oracle determines during compilation if a semantic join should be converted into a classification task and which input dataset contains the labels for the classification. Subsequently, the transformation is performed as a regular rewrite operation by the compiler. The rewritten query is equivalent to the original user query but adds additional operations to: a) reduce the number of input labels to each AI_CLASSIFY call, and b) handle cases where the input of each AI_CLASSIFY call fits in the context window of the user-specified LLM model.
@@ 5.4. AI Aggregation
Earlier, we described an incremental fold strategy for combining intermediate states within a Reduce operation for both and AI_SUMMARIZE_AGG. However, invoking three separate LLM APIs for LLM.Extract, LLM.Combine, and LLM.Summarize introduces significant overhead for small datasets that fit within the model context window. Therefore, a simple ”short-circuit” alternative provides opportunities for throughput improvements by identifying these scenarios and skipping unnecessary incremental fold steps. This simple algorithm change measured an 86.1% latency reduction in AI_SUMMARIZE_AGG on queries using small datasets.
@@ 6. Experimental Evaluation
In this section, we present the experimental evaluation of selected techniques introduced in the preceding sections. Due to space constraints, we focus on a representative subset of results that exemplify the key findings and highlight the practical effectiveness of the proposed methods. All experiments were performed using a production-release version of Snowflake.
@@ 6.1. AI-aware Query Optimization
In this section, we evaluate the performance impact of two compiler optimizations: a) predicate reordering and b) LLM cost-optimized predicate placement with respect to joins. Figure 9 shows the performance impact of predicate reordering for an AISQL query that is shown below. The SQL statement has two predicates in the WHERE clause, an IN predicate and an AI_FILTER predicate on a VARCHAR column. We used a dataset with 1000 articles from the New York Times111https://www.nytimes.com/.
In this experiment, we vary the selectivity of the IN predicate between and , with indicating that all input rows satisfy the IN predicate. We then measure the speedup we get by carefully reordering the predicate evaluation so that the AI_FILTER predicate is evaluated last. As can be seen from the normalized results in Figure 9, applying the AI_FILTER last yields up to a 7 speedup in our experiment. In practice, the actual speedup will depend on several factors, such as the relative costs of the different predicates and their selectivities.
@@
For this experiment, we adjust the ratio of input rows from the left input (NY_ARTICLES_V1), where the AI_FILTER is applied, to the total number of output rows generated by the join. We varied the ratio between and , where means that the join produces twice as many rows as the cardinality of the left input. We compare the optimization that considers the total LLM cost when deciding the placement of AI_FILTER against two other approaches: a) Always Pull-up and, b) Always Push-down. Always Pull-up always ”pulls” AI predicates on top of joins, whereas Always Push-down is the default behavior in Snowflake’s optimizer that always ”pushes” predicates below joins.
As can be seen from the results in Figure 10, neither Always Pull-up nor Always Push-down is optimal across the different output/input ratios. When the ratio is (i.e., a selective join), Always Pull-up performs better. In contrast, Always Push-down is better when the output/input ratio is (i.e., explosive joins). Considering the total cost of LLM when deciding the placement of AI_FILTER predicates (termed AI-aware Optimization) achieves the best performance across the entire range of measured output/input ratios. In practice, the effectiveness of this technique relies on the compiler’s ability to accurately estimate join selectivity, which is not a trivial problem, especially in complicated queries with a large number of joins. Future efforts will explore adaptive optimization techniques, which involve evaluating various execution plans during compilation while collecting runtime statistics, such as the cost and selectivity of AI operators, from sample data to improve decisions regarding AI operator placement.
@@ 6.2. Adaptive Model Cascades
We evaluate the performance impact of model cascades in AI_FILTER operations across seven Boolean classification datasets. As detailed in Section 5.2, the cascading approach uses adaptive threshold learning to route predictions between a lightweight proxy model and a powerful oracle model. We compare three configurations: (1) the baseline approach using only the oracle model Llama3.3-70B for all predictions, (2) the cascade approach combining both models with adaptive threshold learning, and (3) using only the proxy model Llama3.1-8B for all predictions.
Our evaluation methodology executes each query five times to ensure statistical reliability. We measure both execution efficiency (query time) and prediction quality using standard classification metrics: accuracy, precision, recall, and F1 score. All quality assessments use ground-truth labels from the evaluation datasets as a reference.
@@
Performance gains. The cascade approach achieves substantial speedup with mean execution time reduced by 65.5% (from 975.9 to 336.4 seconds), corresponding to a 2.9 speedup. This improvement stems from the cascade’s ability to route the majority of predictions through the lightweight proxy model, reserving the more expensive oracle model only for uncertain cases. For comparison, the proxy model alone runs in 296.2 seconds (3.3 speedup), as expected from its smaller model size, though this comes at a substantial quality cost.
@@
Quality analysis. The cascade maintains strong prediction quality with only modest mean degradation. The F1 score decreases by 4.3% (from 0.812 to 0.777), with balanced precision and recall remaining at 0.784 and 0.794, respectively. This demonstrates that the threshold learning logic effectively identifies rows where the proxy model’s confidence scores are reliable. In contrast, using the proxy model exclusively results in substantially degraded quality, with F1 dropping by 18.8% to 0.659, demonstrating the critical role of the oracle model in handling predictions that fall within the uncertainty region.
As shown in Figure 11, the impact on quality scores varies across datasets. NQ maintains near-identical F1 scores (), indicating that most predictions have clear confidence patterns suitable for routing. IMDB and SST2 show moderate degradation ( and respectively), while BOOLQ and QUORA experience larger drops ( and ). These variations suggest that cascade effectiveness depends on dataset characteristics, i.e., tasks where the proxy model’s uncertainty correlates well with prediction difficulty benefit most from the adaptive routing strategy.
@@
Beyond the speed-quality trade-off, the cascade system provides two additional control mechanisms. Users can restrict the oracle budget by specifying a maximum number of oracle invocations, ensuring predictable costs regardless of data characteristics. The system also supports precision and recall targets, allowing users to specify desired quality thresholds. However, achieving these targets may require higher delegation rates to the oracle model, which reduces the speedup gains.
We also found the cascade algorithm to be resilient to both the degree of parallelism in the SQL execution engine and the number of rows processed by individual workers.
@@ 6.3. Query Rewriting for Semantic Joins
We evaluate AI_CLASSIFY rewrite optimization using eight semantic join benchmarks that cover entity matching (ABTBUY, NASDAQ), document categorization (ARXIV, EURLEX, NYT, CNN), news classification (AG NEWS), and biomedical concept linking (BIODEX). Each benchmark consists of two tables with natural language content, where the join task is to identify semantically related row pairs. For example, in ABTBUY the predicate matches product descriptions from different e-commerce sites that refer to the same item, while in NYT it links news articles to their topical categories. The left table (denoted as in Table 4) and the right table () span 50 to 500 rows, yielding up to 250,000 candidate pairs that must be considered by the naive join approach using filtering.
@@
Table 3 summarizes the mean performance across all benchmarks. Table 4 reports the number of LLM calls, the total execution time (seconds), and standard classification metrics (precision, recall, F1) for both approaches across datasets of varying cardinalities.
Table 4 reports the number of LLM calls, the total execution time (seconds), and standard classification metrics (precision, recall, F1) for both approaches across datasets of varying cardinalities. Performance gains. The query rewriting achieves dramatic speedups across all benchmarks, ranging from (ARXIV) to (CNN), with a mean speedup of . These improvements stem directly from reducing LLM invocations: the baseline requires calls, while the rewrite needs only classifications. For the CNN dataset with 500 rows per table, this reduces 250,000 binary classifications to just 1,000 multi-label classifications, cutting execution time from 4.4 hours to 3.8 minutes. The speedup magnitude correlates strongly with dataset scale; larger Cartesian products yield proportionally greater benefits.
@@
Quality analysis. The impact of rewrite optimization on prediction quality varies by dataset characteristics. On entity matching tasks with clear semantic signals (ABTBUY), both approaches achieve nearly identical F1 scores ( vs ), demonstrating that the rewrite preserves quality when matches are unambiguous. More remarkably, the rewrite substantially improves quality on datasets where the baseline suffers from poor precision. For NASDAQ, the extremely low precision of the baseline () results in an F1 score of , while the rewrite achieves a precision of and an F1 score of . We attribute this to the multi-label classification paradigm: by presenting all candidate labels simultaneously, AI_CLASSIFY enables better comparative reasoning than isolated binary decisions in AI_FILTER. The mean F1 improves by (from to ), with precision increasing from to . Overall, the mean F1 improves by (from to ), with precision increasing from to .

### v3

@@ 5.2. Adaptive Model Cascades
AI operators such as AI_FILTER must often process millions of rows in production workloads. Invoking an LLM on every row introduces prohibitive costs and latency. The AISQL engine addresses this with model cascades: a lightweight proxy model (e.g., Llama3.1-8B) processes all rows and produces a confidence score for each, while a powerful oracle model (e.g., Llama3.3-70B) handles only cases where the proxy is uncertain.
Confidence Scores. For each row , the proxy model produces a score representing the estimated probability that the predicate is satisfied. The score is derived from the model’s output logits: for an AI_FILTER predicate, the proxy generates a binary (yes/no) response and is the softmax probability assigned to the positive-class token.
Two-Threshold Routing. Two learned thresholds and partition rows into three regions based on their proxy scores:
Reject (): predict negative without oracle evaluation.
Accept (): predict positive without oracle evaluation.
Uncertainty (): route to the oracle for reliable classification.
By classifying high- and low-confidence rows with the proxy alone, the cascade reserves oracle invocations for the uncertainty region and the importance sample used for threshold learning.
Adaptive Threshold Learning. AISQL deploys the SUPG-IT algorithm (Liskowski and Schmaus, 2026), which extends the SUPG statistical framework (Kang et al., 2020) to streaming execution with joint precision-recall guarantees. Data is partitioned across parallel workers, each processing its batches independently without inter-worker communication. Within each batch, the algorithm samples a budget fraction of rows from the full batch for oracle labeling using importance sampling with weights proportional to , combined with uniform mixing for coverage. The accumulated oracle labels drive iterative threshold refinement: is set from a weighted ROC curve with a sampling-corrected recall target, while is the minimum threshold whose statistical lower bound on precision meets the precision target. As oracle samples accumulate across batches, confidence bounds tighten, the uncertainty region narrows, and fewer rows require oracle evaluation. Rows remaining in the uncertainty region are routed to the oracle if budget permits; otherwise, the proxy prediction serves as a fallback.
In production, the cascade operates transparently: users issue standard AI_FILTER queries and the engine automatically activates proxy-oracle routing. Users can optionally specify an oracle budget fraction and precision or recall targets to control the cost-quality tradeoff. After each query, the system reports the observed delegation rate so users can adjust these parameters iteratively. Section 6.2 evaluates the cascade’s speed-quality tradeoff across six benchmarks in Snowflake’s production environment. A detailed treatment of SUPG-IT and a complementary calibration-based cascade algorithm, including formal guarantees and convergence analysis, is presented in (Liskowski and Schmaus, 2026).
@@ 5.3. Query Rewriting for Semantic Joins
Semantic joins using AI_FILTER enable natural language-based table joining. For example, let us consider a scenario where a user has two Snowflake tables: (1) “Reviews” contains product reviews, and (2) “Categories” includes names of product categories. A simplified schema of these two tables, along with some sample data, is shown in Figure 8.
@@
Due to the semantic nature of the join condition, efficient join algorithms such as hash-join or sort-merge join, which rely on equality predicates, are not applicable. The AISQL execution engine must therefore treat the semantic join as a cross join, thereby generating all possible pairs from Reviews and Categories and evaluating AI_FILTER for each pair. Therefore, such queries can produce thousands of LLM calls and high inference costs, even for small datasets. More specifically, the above approach requires LLM calls, where and denote the cardinalities of the left and right tables, respectively.
After analyzing a wide range of customer queries, we discovered that, in many cases, semantic joins are equivalent to a multi-label classification problem. For instance, in the previous example, rather than applying AI_FILTER for every review-category pair, we can use the Snowflake AI_CLASSIFY function, which can classify a text or image input into one or more user-specified labels. By re-framing the semantic join as a classification problem, we can instead execute AI_CLASSIFY for every row of table Reviews using the values of column Categories.label as labels, thereby reducing the number of AI inference calls from 24 (4 reviews 6 labels) to just 4 (one for each review), a 6 improvement in the example above.
Automatically detecting such cases is not a trivial task. During the compilation of AISQL queries, we introduced an AI-based oracle that examines every semantic join to determine if it can be rewritten as a multi-label classification problem. The oracle analyzes the user-specified natural language prompt, schema metadata (e.g., column and table names), statistics (e.g., number of distinct values), as well as sample values from each input source. Using all that information, it determines if a semantic join should be converted into a classification task and which input dataset contains the labels for the classification. Subsequently, the transformation is performed as a regular rewrite operation by the compiler. The rewritten query is equivalent to the original user query but adds additional operations to: a) reduce the number of input labels to each AI_CLASSIFY call, and b) handle cases where the input of each AI_CLASSIFY call fits in the context window of the user-specified LLM model. Section 6.3 evaluates the rewrite on eight benchmarks.
@@ 5.4. AI Aggregation
Earlier, we described an incremental fold strategy for combining intermediate states within a Reduce operation for both and AI_SUMMARIZE_AGG. However, invoking three separate LLM APIs for LLM.Extract, LLM.Combine, and LLM.Summarize introduces significant overhead for small datasets that fit within the model context window. Therefore, a simple ”short-circuit” alternative improves throughput by identifying these scenarios and skipping unnecessary incremental fold steps. The optimization achieved an 86.1% latency reduction in AI_SUMMARIZE_AGG on queries using small datasets.
@@ 6. Experimental Evaluation
All experiments run on a production-release version of Snowflake. We evaluate three optimization techniques on public benchmarks: AI-aware query optimization (Section 6.1), adaptive model cascades (Section 6.2), and query rewriting for semantic joins (Section 6.3). For each technique, we measure execution time and prediction quality against baseline approaches.
@@ 6.1. AI-aware Query Optimization
In this section, we evaluate the performance impact of two compiler optimizations: a) predicate reordering and b) LLM cost-optimized predicate placement with respect to joins. Figure 9 shows the performance impact of predicate reordering for an AISQL query shown below. The SQL statement has two predicates in the WHERE clause, an IN predicate and an AI_FILTER predicate on a VARCHAR column. We used a dataset with 1000 articles from the New York Times111https://www.nytimes.com/.
In this experiment, we vary the selectivity of the IN predicate between and , with indicating that all input rows satisfy the IN predicate. We then measure the speedup from reordering the predicate evaluation so that the AI_FILTER predicate is evaluated last. The normalized results in Figure 9 show that applying the AI_FILTER last yields up to a 7 speedup in our experiment. In practice, the actual speedup will depend on several factors, such as the relative costs of the different predicates and their selectivities.
@@
For this experiment, we adjust the ratio of input rows from the left input (NY_ARTICLES_V1), where the AI_FILTER is applied, to the total number of output rows generated by the join. We vary the ratio between and , where means that the join produces twice as many rows as the cardinality of the left input. We compare the optimization that considers the total LLM cost when deciding the placement of AI_FILTER against two other approaches: a) Always Pull-up and, b) Always Push-down. Always Pull-up always ”pulls” AI predicates on top of joins, whereas Always Push-down is the default behavior in Snowflake’s optimizer that always ”pushes” predicates below joins.
Figure 10 shows that neither Always Pull-up nor Always Push-down is optimal across the different output/input ratios. When the ratio is (i.e., a selective join), Always Pull-up performs better. In contrast, Always Push-down is better when the output/input ratio is (i.e., explosive joins). Considering the total cost of LLM when deciding the placement of AI_FILTER predicates (termed AI-aware Optimization) achieves the best performance across the entire range of measured output/input ratios. In practice, the effectiveness of this technique relies on the compiler’s ability to accurately estimate join selectivity, which is not a trivial problem, especially in complex queries with many joins. Future efforts will explore adaptive optimization techniques, which involve evaluating various execution plans during compilation while collecting runtime statistics, such as the cost and selectivity of AI operators, from sample data to improve decisions regarding AI operator placement.
@@ 6.2. Adaptive Model Cascades
We evaluate the performance impact of model cascades in AI_FILTER operations across six Boolean classification datasets. The six datasets are public NLP benchmarks available on HuggingFace, spanning question answering (NQ, BOOLQ), sentiment analysis (IMDB, SST2), duplicate question detection (QUORA), and news veracity classification (FARL). Dataset sizes range from approximately 3,500 to 400,000 rows. As detailed in Section 5.2, the cascading approach uses adaptive threshold learning to route predictions between a lightweight proxy model and a powerful oracle model. We compare three configurations: (1) the baseline approach using only the oracle model Llama3.3-70B for all predictions, (2) the cascade approach combining both models with adaptive threshold learning, and (3) using only the proxy model Llama3.1-8B for all predictions.
We execute each query five times and report the mean. We measure both execution efficiency (query time) and prediction quality using standard classification metrics: accuracy, precision, recall, and F1 score. All quality assessments use ground-truth labels from the evaluation datasets as a reference.
@@
Performance gains. The cascade approach achieves substantial speedup with mean execution time reduced by 65.5% (from 975.9 to 336.4 seconds), corresponding to a 2.9 speedup. The improvement stems from routing the majority of predictions through the lightweight proxy model, reserving the oracle model only for uncertain cases. For comparison, the proxy model alone runs in 296.2 seconds (3.3 speedup), as expected from its smaller model size, though at a significant quality cost.
@@
Quality analysis. The cascade maintains strong prediction quality with only modest mean degradation. The F1 score decreases by 4.3% (from 0.812 to 0.777), with balanced precision and recall remaining at 0.784 and 0.794, respectively. The result confirms that the threshold learning logic identifies rows where the proxy model’s confidence scores are reliable. In contrast, using the proxy model exclusively degrades quality considerably, with F1 dropping by 18.8% to 0.659, underscoring that the oracle model is essential for handling predictions in the uncertainty region.
As shown in Figure 11, the impact on quality scores varies across datasets. NQ maintains near-identical F1 (), while BOOLQ experiences the largest drop (). IMDB, SST2, and QUORA fall between these extremes. The variations suggest that cascade effectiveness depends on dataset characteristics: tasks where the proxy model’s uncertainty correlates well with prediction difficulty benefit most from the adaptive routing strategy.
@@ 6.3. Query Rewriting for Semantic Joins
We evaluate AI_CLASSIFY rewrite optimization using eight semantic join benchmarks, all publicly available on HuggingFace, that cover entity matching (ABTBUY, NASDAQ), document categorization (ARXIV, EURLEX, NYT, CNN), news classification (AG NEWS), and biomedical concept linking (BIODEX). Each benchmark consists of two tables with natural language content, where the join task is to identify semantically related row pairs. For example, in ABTBUY the predicate matches product descriptions from different e-commerce sites that refer to the same item, while in NYT it links news articles to their topical categories. The left table (denoted as in Table 4) and the right table () span 50 to 500 rows, yielding up to 250,000 candidate pairs that must be considered by the naive join approach using filtering. Table 3 summarizes the mean performance across all benchmarks.
@@
Performance gains. The query rewriting achieves large speedups across all benchmarks, ranging from (ARXIV) to (CNN), with a mean speedup of . These improvements stem directly from reducing LLM invocations: the baseline requires calls, while the rewrite needs only classifications. For the CNN dataset with 500 rows per table, this reduces 250,000 binary classifications to just 1,000 multi-label classifications, cutting execution time from 4.4 hours to 3.8 minutes. The speedup magnitude correlates strongly with dataset scale; larger Cartesian products yield proportionally greater benefits.
@@
Quality analysis. The impact of rewrite optimization on prediction quality varies by dataset characteristics. On entity matching tasks with clear semantic signals (ABTBUY), both approaches achieve near-identical F1 scores (), demonstrating that the rewrite preserves quality when matches are unambiguous. The rewrite also substantially improves quality on datasets where the baseline suffers from poor precision. For NASDAQ, the extremely low precision of the baseline () results in an F1 score of , while the rewrite achieves a precision of and an F1 score of . We attribute this to the multi-label classification paradigm: by presenting all candidate labels simultaneously, AI_CLASSIFY enables better comparative reasoning than isolated binary decisions in AI_FILTER. The mean F1 improves by (from to ), with precision increasing from to .